2- an electric bulb is rated 220v and 100 w when it is operated on 110 v the power consumed will be ............. 3-a man with a box on his head is climbing up a ladder the work done by the man on the box is ..... the ans given in my book is positive but i cant understand explain in details




Grade:12

1 Answers

Nikhil Cherukupalli
12 years ago

2) P= V2/R ... Given Pinitial = 100W Vinitial = 220V ...Putting these values in the above equation,R = 222 ohms

    now, Vfinal = 110V.. hence,Pfinal = 110 x 110/ 22x 22 = 25 W

 

3) The general convention in case of Work is; if work is done by the system,it is taken as positive while work done on the system is taken as negetive... Since the work is done by the man on the box, the answers been given as +ve